# Atomic Josephson Parametric Amplifier

### v1

@@ Abstract
Parametric amplification is a fundamental concept in nonlinear physics: When the system’s natural frequency $\omega_{0}$ is periodically modulated in time with frequency $\omega_{\mathrm{d}}$ (the pump), a nearly resonant condition can be achieved for $\omega_{\mathrm{d}}\approx 2\omega_{0}$; in this case, an energy transfer from the pump to the system is induced [1]. This idea was implemented for classical electronics [Roer1994] and then extended to quantum optics [2, 3], semiconductors [4] and superconducting materials [5, 6, 7, 8, 9]. Parametric amplifiers are extensively used in important technological sectors as radio astronomy [10], communications [11], radar systems [12], spectroscopy [13] and sensing [14]. In quantum technology, parametric amplifiers have become relevant for many applications ranging from the generation of squeezed states [5, 6, 15], entanglement and nonclassical states [16, 17], and quantum-limited amplification [18] to quantum computing [19], quantum sensing and metrology [20, 21].
@@
To quantify the overall amplification, we define a dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$, where $S_{\omega}(A_{\mathrm{d}})$ represents the signal peak value and $S_{\omega}(0)$ corresponds to the result without pump. In Fig. 4(a) we demonstrate that $G_{\mathrm{s}}$ initially increases approximately quadratically with $A_{\mathrm{d}}$, and then reaches a maximum for a value of $A_{\mathrm{d}}$, that decreases with increasing $\tilde{I}_{1}=I_{1}/I_{c}$. For larger values of $A_{\mathrm{d}}$, the gain $G_{\mathrm{s}}$ decreases. While the nonlinear increase is well captured by the circuit model, the model fails to describe the decrease beyond the maximum, primarily visible for larger $\tilde{I}_{1}$. Fig. 4(b) presents the signal-to-noise ratio (SNR), defined as $\bar{S}_{\omega}/\Delta\bar{S}_{\omega}$, where we compute the mean signal value $\bar{S}_{\omega}=\bigl{(}S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)\bigr{)}/N_{s}$ and the noise $\Delta\bar{S}_{\omega}$ from the variance determined by $\Delta\bar{S}_{\omega}^{2}=\bigl{(}S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)-\bar{S}_{\omega}\bigr{)}^{2}/N_{s}$. We use $N_{s}=51$, which is the total number of samples in the initial ensemble. The SNR depends on both $A_{\mathrm{d}}$ and $\tilde{I}_{1}$, with an overall value higher than $1$, which confirms robust amplification against the signal noise. This is further corroborated by the enhancement of the signal peak at varying $A_{\mathrm{d}}$ in Fig. 4(c).

### v2

@@ Abstract
Parametric amplification is a fundamental concept in nonlinear physics: When the system’s natural frequency $\omega_{0}$ is periodically modulated in time with frequency $\omega_{\mathrm{d}}$ (the pump), a nearly resonant condition can be achieved for $\omega_{\mathrm{d}}\approx 2\omega_{0}$; in this case, an energy transfer from the pump to the system is induced [1]. This idea was implemented for classical electronics [2] and then extended to quantum optics [3, 4], semiconductors [5] and superconducting materials [6, 7, 8, 9, 10]. Parametric amplifiers are extensively used in important technological sectors as radio astronomy [11], communications [12], radar systems [13], spectroscopy [14] and sensing [15]. In quantum technology, parametric amplifiers have become relevant for many applications ranging from the generation of squeezed states [6, 7, 16], entanglement and nonclassical states [17, 18], and quantum-limited amplification [19] to quantum computing [20], quantum sensing and metrology [21, 22].
@@
To quantify the overall amplification, we define a dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$, where $S_{\omega}(A_{\mathrm{d}})$ represents the signal peak value and $S_{\omega}(0)$ corresponds to the result without pump. In Fig. 4(a) we demonstrate that $G_{\mathrm{s}}$ initially increases approximately quadratically with $A_{\mathrm{d}}$, and then reaches a maximum for a value of $A_{\mathrm{d}}$, that decreases with increasing $\tilde{I}_{1}=I_{1}/I_{c}$. The gain $G_{\mathrm{s}}$ is exemplified in Fig. 4(c) in frequency space. For larger values of $A_{\mathrm{d}}$, the gain $G_{\mathrm{s}}$ decreases. While the nonlinear increase is well captured by the circuit model, the model fails to describe the decrease beyond the maximum, primarily visible for larger $\tilde{I}_{1}$. Fig. 4(b) presents the signal-to-noise ratio (SNR), defined as $\bar{S}_{\omega}/\Delta\bar{S}_{\omega}$, where we compute the mean signal value $\bar{S}_{\omega}=\left\langle S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)\right\rangle$ and the noise $\Delta\bar{S}_{\omega}$ from the variance determined by $\Delta\bar{S}_{\omega}^{2}=\left\langle(S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)-\bar{S}_{\omega})^{2}\right\rangle$. Here, .$\langle$.$\rangle$ . . ⟩ denotes a statistical average over $100$ samples of the ensemble. The SNR depends on both $A_{\mathrm{d}}$ and $\tilde{I}_{1}$, with a visible increase with increasing $A_{\mathrm{d}}$. The maximal regime of the SNR is in the same regime as the maximal gain $G_{\mathrm{s}}$, suggesting that this is the optimal regime of operation.